OpenStudy (anonymous):

Which number is divisible by 3? 1,438 275 3,842 7,221

OpenStudy (anonymous):

If you add the digits of a number and if that sum is divisible by 3, then the number is also divisible by 3.

OpenStudy (anonymous):

7221

OpenStudy (anonymous):

It would have to be 7,221 because when you divide that number by 3 you get 2,407 which is a whole number while the others come out in decimal form

OpenStudy (anonymous):

Add the digits up for each. For example, 1 + 4 + 3 + 8 = 16. Is 16 divisible by 3? 2 + 7 + 5 = 14. Is 14 divisible by 3? 3 + 8 + 4 + 2 = 17. Is 17 divisble by 3? 7 + 2 + 2 + 1 = 12. Is 12 divisble by 3?